The Mach-Zehnder on the PSL table was removed.
A path for 166 MHz modulation in the Mach-Zehnder (MZ) was completely removed, the setup for another path remains the same as before.
Also the photo detector and the CCD for the PMC transmittion were moved to behind the PZT mirror of PMC.
Before removing them, we put an aperture in front of the PD for MC REFL so that we can recover the alignment toward MC by using the aperture.
After the removal we tried to re-align the EOM which imposes the sideband of 29MHz for MC.
We eventually got good alignment of 97% transmissivity at the EOM ( the power of the incident beam is 1.193W and trans was 1.160W )
And then we aligned the beam going to MC by guiding the reflected beam to the aperture we put. This was done by using the steering mirrors on the periscope on the corner of the PSL table.
Now MC got locked and is successfully resonating with TEM00.